OpenStudy (anonymous):

Solve A = 5(b + 2c) for c in terms of A and b. You must show all work.

OpenStudy (anonymous):

A = 5(b + 2c) for c A = 5b + 10c ->A - 5b = 10c Thus c = A/10 - b/2

OpenStudy (anonymous):

First you want to simplify as much as possible by distributing the 5: A = 5(b + 2c) A = 5b + 10c Next you want to isolate the 10c, so we need to get get rid of the 5b, but what youd o to one side you have to do to the other so A - 5b = 5b + 10c - 5b A - 5b = 10c Now you want to get c all by itself, so divide by ten on both sides: (A-5b)/10 = c tada!